# COSMIC: Compress Satellite Images Efficiently via Diffusion Compensation

### v1

@@ 2.1 Earth observation missions on satellites
Earth observation missions (e.g. NASA’s Landsat Program [49]) involve the use of satellite photos to monitor and collect data for tasks like forestry [5], agriculture [14, 41], land degradation [15], land use and land cover [40], biodiversity [34], and water resource [35, 47]. Traditional earth observation missions rely on a pipeline in which satellites take photos and then send them back to ground stations for analysis. Recently, along with the rapid cost of launching rockets and manufacturing nanosatellites [21, 6], the photos on satellites are rapidly increasing which brings novel challenges for transmitting photos back to the ground. A recent promising approach is to deploy embedded GPUs on satellites to support DNN models to either filter useless data before transmission or make partial processing tasks on satellites. For instance, ESA’s satellite Phi-Sat-1 [25] first deploy Intel VPU on satellite to support DNN models for filtering useless photos (e.g. covered by clouds) that can save 30%+ transmission volume. OroraTech has launched AI nanosatellites with the NVIDIA Jetson Xavier NX for wildfire detection [4], and Orbital Sidekick uses NVIDIA Jetson AGX Xavier as the AI engine at the edge of the satellite to detect gas pipeline leaks [2]. However, due to the inelastic computational capabilities of onboard satellites and limited power supply only from the sunshine (i.e., up to 15 Watt for GPUs on satellites [4]), a certain amount of images are still needed to be transmitted back to the ground. This brings an urgent need for satellite-specific image compression methods.
@@ 5.1 Setup
Dataset. We use the function Map of the World (fMoW) [13], which has 62 categories, and in which each image is paired with different types of metadata features, as our training data and test data. For training data, we randomly crop the image to a resolution of $256\times 256$ pixels. For information collected by satellite sensors as metadata, we choose UTM zone, Timestamp, GSD, Cloud cover, Off-nadir Angle, Target Azimuth, Sun Azimuth, and Sun Elevation provided by fMoW.
@@ 5.2 Comparison with baselines
RD performance. Figure 3 shows the comparison results with baselines on two test sets. The dotted line in the figure represents the baselines, and the solid line represents COSMIC. We demonstrate results from two perspectives, i.e., distortion and perception. Across all transmission rates, COSMIC surpasses the baselines in terms of LPIPS, and FID. At low bpp, the MS-SSIM of COSMIC is lower compared to the baseline. This is mainly because as the bpp decreases, the encoder extracts less information, and during the decompression process, there is a greater reliance on diffusion-based generation (more details in Sec. 5.4). Additionally, due to the degraded feature extraction of the lightweight encoder, the feature obtained at low bpp is insufficient to guide the diffusion process in generating high-fidelity images. As the bpp increases, the latent coding contains more features, resulting in a significant improvement in the MS-SSIM of COSMIC, demonstrating SOTA performance. Note that the sensor data is transmitted to the ground by default in earth observation missions. Besides, the volume of these sensor data is negligible compared with images so we do not consider them when counting bpp. We show more results of more metrics and baselines including VAE-based methods in Appendix B which COSMIC achieves the SOTA performance on all 6 perceptual metrics.
@@ 5.4 Ablation study
Figure 6: Ablation study with different variants of COSMIC. “w/o DC” indicates that diffusion compensation was not used during decoding. “w/o ME” denotes that we remove the metadata encoder. To show the effect of the metadata encoder, we mark the finer differences with red boxes in the figure.
@@ 4 $\lfloor\mathbf{y}\rceil=\,\,\mathrm{Quantization}\left(\mathrm{y},\mathrm{\zeta}\right)$;
A.2 Detailed architecture of COSMIC
The overall framework of COSMIC has been illustrated in the main paper. Here, we provide the detailed architecture of image encoder $\tilde{\mathcal{E}}$ and image decoder $\mathcal{D}$ respectively in Table 3.
Table 3: Detailed network structure of $\tilde{\mathcal{E}}$ and $\mathcal{D}$.
Architecture of image encoder $\tilde{\mathcal{E}}$
input referenced image
layer1 Conv($3\times 192\times 5\times 5$,s=2), GND(192)
layer2 Conv($192\times 192\times 5\times 5$,s=2), GND(192)
layer3 Conv($192\times 192\times 5\times 5$,s=1), GND(192)
layer4 Conv($192\times 8\times 5\times 5$,s=1)
output $\mathrm{z}_{0}$
Architecture of image decoder $\mathcal{D}$
input $\lfloor\mathbf{y}\rceil$, $\mathrm{z}_{0}\prime$
layer1 Conv($196\times 192\times 5\times 5$,s=1), IGND(192)
layer2 Conv($192\times 192\times 5\times 5$,s=1), IGND(192)
layer3 Conv($192\times 192\times 5\times 5,$s=2), IGND(192)
layer4 Conv($192\times 3\times 5\times 5$,s=2)
output reconstructed image

### v2

@@ 2.1 Earth observation missions on satellites
Earth observation missions (e.g. NASA’s Landsat Program [49]) involve the use of satellite photos to monitor and collect data for tasks like forestry [5], agriculture [14, 41], land degradation [15], land use and land cover [40], biodiversity [34], and water resource [35, 47]. Traditional earth observation missions rely on a pipeline in which satellites take photos and then send them back to ground stations for analysis. Recently, along with the reduced cost of launching rockets and manufacturing nanosatellites [21, 6], the photos on satellites are rapidly increasing which brings novel challenges for transmitting photos back to the ground. A recent promising approach is to deploy embedded GPUs on satellites to support DNN models to either filter useless data before transmission or make partial processing tasks on satellites. For instance, ESA’s satellite Phi-Sat-1 [25] first deploy Intel VPU on satellite to support DNN models for filtering useless photos (e.g. covered by clouds) that can save 30%+ transmission volume. OroraTech has launched AI nanosatellites with the NVIDIA Jetson Xavier NX for wildfire detection [4], and Orbital Sidekick uses NVIDIA Jetson AGX Xavier as the AI engine at the edge of the satellite to detect gas pipeline leaks [2]. However, due to the inelastic computational capabilities of onboard satellites and limited power supply only from the sunshine (i.e., up to 15 Watt for GPUs on satellites [4]), a certain amount of images are still needed to be transmitted back to the ground. This brings an urgent need for satellite-specific image compression methods.
@@ 5.1 Setup
Dataset. We use the function Map of the World (fMoW) [13], which has 62 categories, and in which each image is paired with different types of metadata features, as our training data and test data. For training data, we randomly crop the image to a resolution of $256\times 256$ pixels. For information collected by satellite sensors as metadata, we choose Coordinates, Timestamp, GSD, Cloud cover, Off-nadir Angle, Target Azimuth, Sun Azimuth, and Sun Elevation provided by fMoW.
@@ 5.2 Comparison with baselines
RD performance. Figure 3 shows the comparison results with baselines on two test sets. The dotted line in the figure represents the baselines, and the solid line represents COSMIC. We demonstrate results from two perspectives, i.e., distortion and perception. Across all transmission rates, COSMIC surpasses the baselines in terms of LPIPS, and FID. At low bpp, the MS-SSIM of COSMIC is lower compared to the baseline. This is mainly because as the bpp decreases, the encoder extracts less information, and during the decompression process, there is a greater reliance on diffusion-based generation (more details in Sec. 5.4). Additionally, due to the degraded feature extraction of the lightweight encoder, the feature obtained at low bpp is insufficient to guide the diffusion process in generating high-fidelity images. As the bpp increases, the latent coding contains more features, resulting in a significant improvement in the MS-SSIM of COSMIC, demonstrating SOTA performance. Note that the sensor data is transmitted to the ground by default in earth observation missions. Besides, the volume of these sensor data is negligible compared with images so we do not consider them when counting bpp. We show more results of more metrics and baselines including other VAE-based methods in Appendix B which COSMIC achieves the SOTA performance on all 6 perceptual metrics.
@@ 5.4 Ablation study
Figure 6: Ablation study with different variants of COSMIC. “w/o DC” indicates that diffusion compensation is not used during decoding. “w/o CAM”denotes the CAM module is removed in the encoder. “w/o ME” denotes that we remove the metadata encoder.